OpenStudy (anonymous):

simplify: (-3x^-2)(4x^4)

OpenStudy (anonymous):

\[\left( -3x ^{-2} \right)\left( 4x ^{4} \right)\]

jhonyy9 (jhonyy9):

\[(1/(-3x ^{2}) )*(4x ^{4})=(4x ^{4})/(-3x ^{2})=(4x ^{2})/(-3)= -4x ^{2}/3\]